OpenStudy (josedavid):

Choose the equation that could be used to find three consecutive integers whose sum is 75. n + (n + 1) + (n + 2) = 75 n + (n + 2) + (n + 4) = 75 n + (n + 1) + (n + 3) = 75 n + (n - 1) + (n - 3) = 75

OpenStudy (josedavid):

@ameenbriscoe13 @quickstudent @queenoftherings @wio @Werewolfprincess @willsgirl16 @WhiteFangIII @eliassaab @emmigrace222 @elone000 @elone000 @radar @razor99 @roz_mar23 @romanortiz65 @ROSEQUEEN2014 @Richbgsn @terenzreignz @tanya123 @Tazmaniadevil @TheAsker2002 @Thatkid_marc @Tuta @Yuba @YungXxGodXx @yazminva @uri @uChezzy @Ultra_LoveBBY @unicornsandmarshmellows @iambatman @iceicebaby @ilovehim121511 @itsbribro @invinceble @issues101@One098 @oswin @Olaelyamany @OdinMW @openstudysignup @PookieBear555 @pookie0799 @patricksweeney @PrincesssCDE @Prissy416 @animal_lover36 @AngelPSP @appleduardo @Angels @AquaStone316 @alprincenofl @secret66 @Secret-Ninja @sleepyjess

OpenStudy (anonymous):

n+(n+2)+(n+4)= 75

OpenStudy (anonymous):

n+(n+1)+(n+2)=75 Solve for n 3n+3=75 3n=75-3 3n=72 n=24 24+(24+1)+(24+2)=75 24+25+26=75